OpenStudy (anonymous):

Need to form a polynomial function f(x) with a degree of 3; zeros -9, and 3-i. I am super confused!!! I think the problem should look something like this (X+9)(x-3-i)(x+3+i) can someone help me please :-)

OpenStudy (anonymous):

yes, close

OpenStudy (anonymous):

it should be \[(x+9)(x-(3-i))(x-(3+i))\] and your first job is to multiply out \[(x-(3-i))(x-(3+i))\] which is not nearly as hard as it looks

OpenStudy (anonymous):

we can use one of three methods to find that quadratic, the one with zeros \(3+i\) and \(3+i\) one method is to work backwards \[x=3+i\] \[x-3=i\] \[(x-3)^2=-1\] \[x^2-6x+9=-1\] \[x^2-6x+10=0\] so it is \[x^2-6x+10\]

OpenStudy (anonymous):

another it to recall that if the zero is \(a+bi\) the quadratic is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

in your case \(a=3,b=1\) so you get \[x^2-2\times 3x+3^2+1^2=x^2-6x+10\]